An urn contains ten marbles, of which five are green, two are blue, and three are red. Three marbles are to be drawn from the urn, one at a time without replacement. What is the probability that all three marbles drawn will be green?
Answer by galactus(183)

The are 10 marble altogether. The prob. of choosing a green on the first draw is 5/10=1/2. Since we do not put it back, there are 9 marbles, 4 of which are green. The second green has prob. 4/9
The third, 3/8
So, the prob. all three are green is (1/2)(4/9)(3/8)=1/12.